OpenStudy (anonymous):

Reduce the fraction 20/100.

OpenStudy (anonymous):

i think is 0.02

OpenStudy (anonymous):

answer is 1/5 b'coz we are reducing the given fraction.

OpenStudy (anonymous):

sorry forget about about that true if we do this \[20/100 \div20/20\]

OpenStudy (anonymous):

Firstly, note that 10 is a common factor of both numbers, so we can write the fraction as: \[ \frac{2(10)}{2(50)} \]The twos cancel out, leaving us with: \[ \frac{10}{50} \]Notice how 10 is a common factor, so we have: \[ \frac{1}{5} \]

OpenStudy (anonymous):

The decimal number 0.75 means:

OpenStudy (anonymous):

the decimal number 0.75 means what? a)there are seventy-five parts of a whole unit that has 100 parts or b) seventy-five tenths of a whole unit or c) that a whole unit consist of sevrnty-five parts or d) there are seven-and-a-half parts of a whole number.